OpenStudy (anonymous):

Please help I don't understand this one at all. Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 3, – 25), and x intercepts at x = -8 and x = 2. y = x2 + 6x -_____

OpenStudy (anonymous):

graph x2 + 6x see where vertex is at use google or your graph calc

OpenStudy (anonymous):

x^2 + 6*x

OpenStudy (anonymous):

the vertex is -3,-9 you dont have to do nothing to the -3 but you do have to shift it down till it reaches -25 so -25=x+(-9) x=-16 so y = x2 + 6x -_____ y = x^2 + 6x -16 :)

OpenStudy (anonymous):

\[y=(x+8)(x-2)\] \[y=x^2+6x-16\] and you can check that if you replace \(x\) by -3 you get \[(-3)^2+6\times -3-16=9-18-16=-9-16=-25\]

OpenStudy (anonymous):

dude that is too much y = x2 + 6x you are given it ask for y = x2 + 6x - c All you have to do is shift it down :)

OpenStudy (anonymous):

Thank you both!